OpenStudy (anonymous):

Write the partial fraction decomposition of the rational expression. Check your result algebraically. (x^2+6x+3)/(x^3+x)

OpenStudy (mathmale):

Let's break up your expression into partial fractions. Noting that the denominator factors into \[x(x^2+1)\]

OpenStudy (anonymous):

explain please? i have exam tomorrow

OpenStudy (anonymous):

yes i got that part

OpenStudy (mathmale):

We get \[\frac{ x^2+6x+3 }{ x^3+x }=\frac{ A }{ x }+\frac{ Bx+C }{ x^2+1 }\]

OpenStudy (anonymous):

yes!

OpenStudy (mathmale):

Before we move on, please ask any questions that you may have.

OpenStudy (anonymous):

i understand so far! thank you for helping

OpenStudy (mathmale):

Good! Now we'd like to eliminate the fractions. Look at the "A" term. Missing from its denominator is (x^2+1), RIGHT? so, multiply both A and x by (x^2+1). OK?

OpenStudy (anonymous):

ummmmmmmmmmmmm

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

i get it, multiply A and x by (x^2+1) and x with the other term

OpenStudy (anonymous):

sorry if i am wrong

OpenStudy (mathmale):

Just for verification, we need to obtain the following:\[\frac{ A(x^2+1) }{ x(x^2+1) }\]

OpenStudy (mathmale):

Can you now do the B term yourself? Mult. num. and denom. each by x.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wouldn't be x^3 +x then for denominator?

OpenStudy (anonymous):

after multiplying by x

OpenStudy (mathmale):

Yes, but there's no point in doing this multiplication. We're trying to get rid of the denom.

OpenStudy (anonymous):

oh so it should be (Bx+C)(x)

OpenStudy (mathmale):

As a matter of fact, we should now have \[x^2+6x+3=A(x^2+1)+Bx\]

OpenStudy (mathmale):

with no denominators. Can you believe that? Yes, sorry, x(Bx +C).

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

i see where i went wrong but i am on track

OpenStudy (mathmale):

Good. Finish this on your own and let me know if and when you have further questions. Glad to see you again on OpenStudy.

OpenStudy (anonymous):

hey thanks

OpenStudy (mathmale):

My pleasure!